A straight line L passes though the point (0, -1) and is parallel to the straight line y=3x+1.
i. Write down the equation of the straight line L.
ii. Given that L passes through the point (2,b) find the value of b.

Found 2 solutions by Alan3354, RAY100:
Answer by Alan3354(69443)

A straight line L passes though the point (0,-1) and is parallel to the straight line y=3x+1.
-------------
The slope, m, of the line y=3x+1 is 3.
------------
i. Write down the equation of the straight line L.
Use (always):
y-y1 = m*(x-x1) (x1,y1) is the point (0,-1)
y+1 = 3x
y = 3x-1 is the line L.
ii. Given that L passes through the point (2,b) find the value of b.
y = 3x-1
x = 2
y = 5 = b

Answer by RAY100(1637)

Remember y= mx +b
.
y=3x +1,,,,,,m=3,,,,,parallel to this also has m=3
.
parallel line,,,,y= 3x +b,,,thru (0,-1),,,,subst
.
(-1) = 3(0) +b
.
b= -1
.
or parallel line (L),,,,y= 3x -1 (i)
.
if y=3x-1,,,,thru (2,b),,,,subst
.
(b) = 3(2) -1,,,,,b= 5,,,,,(ii)
.